OpenStudy (anonymous):

log2+log3+logx=log5 findx

OpenStudy (ash2326):

@Patory I'll give you a hint. Apoorv please don't give answer directly

OpenStudy (anonymous):

we have loga+logb=logab now u can solve

OpenStudy (anonymous):

is it ok ash2326??

OpenStudy (ash2326):

There are two log properties that'd be essential here \[\log a+ \log b=\log ab\] and \[log_{10} x=y=> 10^y=x\] Can you solve now?